Dundar Mifflin makes electronic products for the RCMP. The following data is for the first six months:

            

Direct Labor Hours

Manufacturing Overhead

January

45,000

$295,000

February

59,900

$307,575

March

57,000

$323,000

April

58,000

$320,000

May

27,800

$158,600

June

29,000

$152,500

Required:

  1. Use the high-low method to estimate the cost formula                                (3 marks)

  1. Estimate the total overhead cost at an activity level of 432000 machine hours, using the separate estimates you obtained for its components.                                                                         (2 marks)
Answer #1

rate positively ..

Direct Labor Hours Manufacturing Overhead
January 45,000 $295,000
February 59,900 $307,575
March 57,000 $323,000
April 58,000 $320,000
May 27,800 $158,600
June 29,000 $152,500
Maximum 59,900 323,000
minimum 27,800 152,500
Delta 32,100 170,500
Variable cost= $               5.31
170500/32100
Fixed cost=323000-59900*5.31 $             4,840
Therefore cost equation = 4840+5.31*x
4840+5.31*x
total overhead cost at an activity level of 432000=
432000*5.31+4840= $      2,299,419
Ans= $      2,299,419